# Neural reproducing kernel Banach spaces and representer theorems for deep networks

### v1

@@ 2. Preliminaries
where the sum converges absolutely in the $\|\cdot\|_{\mathcal{Y}}$ norm. We denote by $\mathcal{M}(\Theta,\mathcal{Y})$ the space of vector measures on $\Theta$ with values in $\mathcal{Y}$. If $\mu\in\mathcal{M}(\Theta,\mathcal{Y})$, for all $A\in\mathcal{B}(\Theta)$ we define
$|\mu|(A)=\sup_{\{A_{i}\}}\sum_{i}\|\mu(A_{i})\|_{\mathcal{Y}}\ ,$
where the supremum is taken over all partitions of $A$. Then, $|\mu|$ is a bounded positive measure on $\Theta$, which we call the total variation of $\mu$. The space $\mathcal{M}(\Theta,\mathcal{Y})$ is a Banach space with respect to the total variation norm
$\|\mu\|_{\operatorname{TV}}=|\mu|(\Theta)\ .$
Remark 2.5.
Fix a vector measure $\mu$. Since $\Theta$ is a second countable space, for all $y^{\prime}\in\mathcal{Y}^{\prime}$, ${}_{\mathcal{Y}}\langle\mu(\cdot),y^{\prime}\rangle_{\mathcal{Y}^{\prime}}$ is a scalar regular (Radon) measure, hence $\mu$ is a regular vector measure [21]. Since the elements of $\mathcal{M}(\Theta,\mathcal{Y}^{\prime})$ are regular measures, the space $\mathcal{M}(\Theta,\mathcal{Y}^{\prime})$ can be identified with the dual of $\mathcal{C}_{0}(\Theta,\mathcal{Y})$, the space of continuous functions from $\Theta$ to $\mathcal{Y}$ that vanish at infinity, by (a generalization of) the Riesz representation theorem. In particular, $\mathcal{M}(\Theta,\mathcal{Y}^{\prime})$ can be endowed with the product $*{}^{*}$ topology, with respect to which the closed balls are compact. This result was proved in [23] for compact $\Theta$ (see also [21]), and then extended in [8] to locally compact $\Theta$ in the case where $\mathcal{Y}$ is Hilbert. The proof in [8] holds true with the obvious modifications if $\mathcal{Y}$ is Banach.
3. Neural RKBS
In this section, we introduce particular classes of RKBS, incorporating general properties inspired by neural networks architectures. In particular, we construct deep and neural RKBS, and show that neural networks as intended in 2.1 are elements of neural RKBS. In this way, we provide neural networks with a Banach space functional structure, viewing them as finite representations of an underlying nonparametric model.
Integral RKBS
Let $\Theta,\Xi$ be two locally compact, second countable topological spaces. Let $\mathcal{X}$ be a set, and let $\mathcal{Y}$ be a RKBS on $\Xi$. Let
@@ 3. Neural RKBS
$f_{\mu}:\mathcal{X}\to\mathcal{Y}\qquad f_{\mu}(x)=\phi(x)\mu=\int_{\Theta}\rho(x,\theta){\rm d}\mu(\theta)$ (2)
@@
=$\mathcal{H}=\{f_{\mu}:\mathcal{X}\to\mathcal{Y}:\mu\in\mathcal{M}(\Theta,\mathcal{Y})\}$
@@
The space $\mathcal{H}$ thus defined is a (vector valued) RKBS, which we call an integral RKBS. Moreover, we call $\Theta$ and $\Xi$ the parameter spaces of $\mathcal{H}$, and $\rho$ its basis functions.
Since $\mathcal{Y}$ is a RKBS, for every $\xi\in\Xi$ there is $\delta_{\xi}\in\mathcal{Y}^{\prime}$ such that $y(\xi)={}_{\mathcal{Y}}\langle y,\delta_{\xi}\rangle_{\mathcal{Y}^{\prime}}$ for all $y\in\mathcal{Y}$. Thus, defining the (scalar) measure
$\mu(E\mid\xi)={}_{\mathcal{Y}}\langle\mu(E),\delta_{\xi}\rangle_{\mathcal{Y}^{\prime}}\qquad E\in\mathcal{B}(\Theta)\qquad\xi\in\Xi\ ,$
for all $x\in\mathcal{X}$ we have
$f_{\mu}(x)(\xi)=\int_{\Theta}\rho(x,\theta){\rm d}\mu(\theta\mid\xi)\ .$
@@
Let $L$ be a positive integer. Take a set $\mathcal{X}_{0}$ and Banach spaces $\mathcal{X}_{1},\dots,\mathcal{X}_{L+1}$. For $\ell=0,\dots,L$, take RKBS $\mathcal{H}_{\ell}$ on $\mathcal{X}_{\ell}$ with values in $\mathcal{X}_{\ell+1}$. The direct sum
@@
$\|f\|_{\mathcal{H}}=\|f_{0}\|_{\mathcal{H}_{0}}+\cdots+\|f_{L}\|_{\mathcal{H}_{L}}\qquad f=f_{0}\oplus\cdots\oplus f_{L}\ .$
To every $f=f_{0}\oplus\cdots\oplus f_{L}\in\mathcal{H}$ we assign the function $f^{\text{deep}}:\mathcal{X}_{0}\to\mathcal{X}_{L+1}$ defined by
$f^{\text{deep}}=f_{L}\circ\cdots\circ f_{0}\ .$
The linear space $\mathcal{H}$ parametrizes the nonlinear space
$\mathcal{H}^{\text{deep}}=\{f^{\text{deep}}:\mathcal{X}_{0}\to\mathcal{X}_{L+1}:f\in\mathcal{H}\}\ .$
With slight abuse of language, we call the non-linear space $\mathcal{H}^{\text{deep}}$ a RKBS of depth $L$ parametrized by the RKBS $\mathcal{H}=\mathcal{H}_{0}\oplus\cdots\oplus\mathcal{H}_{L}$. If $L>1$, we refer to $\mathcal{H}^{\text{deep}}$ as a deep RKBSp.
Moreover, we call $\mathcal{X}_{\ell}$ the layer spaces of $\mathcal{H}$. In particular, $\mathcal{X}=\mathcal{X}_{0}$ is the input space, $\mathcal{Y}=\mathcal{X}_{L+1}$ is the output space, and $\mathcal{X}_{\ell}$, $\ell=1,\dots,L$, are the hidden layer spaces.
@@
Let $L$ be a positive integer. If the RKBS $\mathcal{H}_{\ell}$ are integral,
=$\mathcal{H}_{\ell}=\{f_{\mu_{\ell}}:\mathcal{X}_{\ell}\to\mathcal{X}_{\ell+1}:\mu_{\ell}\in\mathcal{M}(\Theta_{\ell},\mathcal{X}_{\ell+1})\}\ ,$
@@
the corresponding nonlinear space
$\mathcal{H}^{\text{deep}}=\{f^{\text{deep}}:\mathcal{X}_{0}\to\mathcal{X}_{L+1}:f^{\text{deep}}=f_{L}\circ\cdots\circ f_{0}\ \ ,f_{\ell}=f_{\mu_{\ell}}\in\mathcal{H}_{\ell}\ ,\ell=0,\ldots,L\}$
@@
is called an integral RKBS of depth $L$ parametrized by $\mathcal{H}=\mathcal{H}_{0}\oplus\cdots\oplus\mathcal{H}_{L}$ and, if $L>1$, a deep integral RKBS.
A function $f:\mathcal{X}_{0}\to\mathcal{X}_{L+1}\in\mathcal{H}^{\text{deep}}$ has the form $f(x)=x^{(L+1)}$ where
@@
$\begin{cases}x^{(0)}=x&\in\mathcal{X}_{0}\\[5.0pt] x^{(\ell+1)}=\displaystyle{\int_{\Theta_{\ell}}\rho_{\ell}(x^{(\ell)},\theta_{\ell}){\rm d}\mu_{\ell}(\theta_{\ell})}&\in\mathcal{X}_{\ell+1}\qquad\ell=0,\dots,L\ .\end{cases}$
Neural RKBS

### v2

@@ 3. Vector measures, integral RKBS and Neural Networks
where the sum converges unconditionally in the $\|\cdot\|_{\mathcal{Y}}$-norm.
Pettis theorem [14, Thm.1 IV.10.1] shows that a set function $\mu\colon\mathcal{B}(\Theta)\to\mathcal{Y}$ is a vector measure if and only if $\mu_{y^{\prime}}={}_{\mathcal{Y}}\langle\mu(\cdot),y^{\prime}\rangle_{\mathcal{Y}^{\prime}}$ is a $\sigma$-additive measure for all $y^{\prime}\in\mathcal{Y}^{\prime}$.
Definition 3.4 (Variation of a vector measure).
Let $\mu$ be a vector measure on $\Theta$ with values in $\mathcal{Y}$. The variation of $\mu$ is the function $|\mu|\colon\mathcal{B}(\Theta)\to[0,+\infty]$ defined by
$|\mu|(A)=\sup_{\{A_{i}\}}\sum_{i}\|\mu(A_{i})\|_{\mathcal{Y}}\qquad A\in\mathcal{B}(\Theta),$
where the supremum is taken over all finite partitions of $A$. If $|\mu|(\Theta)<+\infty$, the measure $\mu$ is called a vector measure of bounded variation.
The space $\mathcal{M}(\Theta,\mathcal{Y})$ of vector measures of bounded variation is a Banach space with respect to the norm
$\|\mu\|_{\rm TV}=|\mu|(\Theta).$
If $\mu\in\mathcal{M}(\Theta,\mathcal{Y})$, its variation $|\mu|$ is a finite positive measure on $\Theta$, see [13, 1.A.10].
The integration of a scalar function $\varphi$ with respect a vector measure of bounded variation can be defined as the Bochner integral of a vector valued function, see [13, Ch.1 Section D] and [14, IV.10]. In particular, a measurable scalar function $\varphi$ is integrable with respect to $\mu$ if and only if $\varphi$ is integrable with respect to $|\mu|$. If $\varphi=\sum_{i}t_{i}{\chi}_{A_{i}}$ is a simple function, then
$\int_{\Theta}\varphi(\theta)d\mu(\theta)=\sum_{i}t_{i}\mu(A_{i})\in\mathcal{Y},$
and the integral of an arbitrary $|\mu|$-integrable functions is defined via the density of simple functions.
If $\mathcal{Y}$ has the Radon-Nikodym property [13, Ch. 1.G], as it happens if $\mathcal{Y}$ is reflexive, then $\mu$ has density with respect to $|\mu|$, i.e. there exists a function $g\colon\Theta\to\mathcal{Y}$ such that $\|g(\theta)\|_{\mathcal{Y}}=1$ for all $\theta\in\Theta$, $g$ is Bocnher integrable with respect to $|\mu|$, and
$\int_{\Theta}\varphi(\theta)d\mu(\theta)=\int_{\Theta}\varphi(\theta)g(\theta)d|\mu|(\theta),$
for all $\mu$-integrable scalar functions $\varphi:\Theta\to\mathbb{R}$. If $\mathcal{Y}$ does not have the Radon-Nikodym property, nevertheless there always exists a function $g:\Theta\to\mathcal{Y}$ such that
  1. i)
    for all $y^{\prime}\in\mathcal{Y}^{\prime}$, $_{\mathcal{Y}}\langle g(\cdot),y^{\prime}\rangle_{\mathcal{Y}^{\prime}}$ is $|\mu|$-integrable;
  2. ii)
    for all $\theta\in\Theta$, $\|g(\theta)\|_{\mathcal{Y}}=1$;
  3. iii)
    for all scalar function $\varphi$ that are $\mu$-integrable and $y^{\prime}\in\mathcal{Y}^{\prime}$
    $_{\mathcal{Y}}\langle\int_{\Theta}\varphi(\theta)d\mu(\theta),y^{\prime}\rangle_{\mathcal{Y}^{\prime}}=\int\varphi(\theta){}_{\mathcal{Y}}\langle g(\theta),y^{\prime}\rangle_{\mathcal{Y}^{\prime}}d|\mu|(\theta),$
see [13, Ch.1 Theorem 34].
Vector integral RKBS and neural networks
We now introduce particular classes of RKBS. In particular, we extend the infinite-width limit (4) of shallow neural networks from scalar to vector-valued functions. This result is of independent interest and will be crucial for the extension from shallow to deep networks.
Definition 3.5 (Integral RKBS).
Let $\Theta$ be a locally compact, second countable topological space, $\mathcal{X}$ a set, and $\mathcal{Y}$ a Banach space. Let
@@
$f_{\mu}:\mathcal{X}\to\mathcal{Y},\qquad f_{\mu}(x)=\phi(x)\mu=\int_{\Theta}\rho(x,\theta){\rm d}\mu(\theta),$ (5)
@@
$\mathcal{H}=\{f_{\mu}:\mathcal{X}\to\mathcal{Y}:\mu\in\mathcal{M}(\Theta,\mathcal{Y})\},$
@@
The space $\mathcal{H}$ thus defined is a (vector-valued) RKBS, which we call an integral RKBS. Moreover, we call $\Theta$ the parameter spaces of $\mathcal{H}$, and $\rho$ its basis function.
Remark 3.6.
If $\rho(x,\cdot)$ is bounded for all $x\in\mathcal{X}$, (5) is still well defined. The stronger assumption that $\rho(x,\cdot)\in\mathcal{C}_{0}(\Theta)$ ensures that the map $f\mapsto f(x)$ is weakly continuous, see Remark˜5.2.
Remark 3.7.
Scalar valued integral RKBS correspond to the choice $\mathcal{Y}=\mathbb{R}$ and coincide with the setting considered in [3].
4. Deep Integral and Neural RKBS
In this section, we start introducing deep integral RKBS. Then, we derive representer theorems on these spaces, which show how optimal solutions minimizing empirical objectives can be taken to have a finite width at every layer.
@@ 4. Deep Integral and Neural RKBS
The first step is to take direct sums of RKBS and define the deep RKBS as a composition of elements in the direct sum.
Definition 4.1 (Deep RKBS).
Let $\mathcal{X}$ be a set and $\mathcal{Y}$ a Banach space. Fix a positive integer $L\geq 1$. Take a set $\mathcal{X}_{0}=\mathcal{X}$ and Banach spaces $\mathcal{X}_{1},\dots,\mathcal{X}_{L+1}=\mathcal{Y}$. For $\ell=0,\dots,L$, take RKBS $\mathcal{H}_{\ell}$ on $\mathcal{X}_{\ell}$ with values in $\mathcal{X}_{\ell+1}$. The direct sum
@@
$\|f\|_{\mathcal{H}}=\|f_{0}\|_{\mathcal{H}_{0}}+\cdots+\|f_{L}\|_{\mathcal{H}_{L}},\qquad f=f_{0}\oplus\cdots\oplus f_{L}.$
To every $f=f_{0}\oplus\cdots\oplus f_{L}\in\mathcal{H}$ we assign the function $f^{\rm deep}:\mathcal{X}\to\mathcal{Y}$ defined by
$f^{\rm deep}=f_{L}\circ\cdots\circ f_{0},$
and we set
$\mathcal{H}^{\rm deep}=\{f^{\rm deep}:\mathcal{X}\to\mathcal{Y}:f\in\mathcal{H}\}\$
endowed with the complexity measure $\Phi:\mathcal{H}^{\rm deep}\to[0,+\infty)$ given by
$\Phi(f^{\rm deep})=\inf\{\|g\|_{\mathcal{H}}:g\in\mathcal{H}\ \text{such that }g^{\rm deep}=f^{\rm deep}\}.$ (7)
With a slight abuse of language, we call the nonlinear space $\mathcal{H}^{\rm deep}$ an RKBS of depth $L$ induced by the Banach space $\mathcal{H}$. If $L>1$, we refer to $\mathcal{H}^{\rm deep}$ as a deep RKBS. Moreover, we call $\mathcal{X}_{\ell}$ the layer spaces of $\mathcal{H}$. In particular, $\mathcal{X}=\mathcal{X}_{0}$ is the input space, $\mathcal{Y}=\mathcal{X}_{L+1}$ is the output space, and $\mathcal{X}_{\ell}$, $\ell=1,\dots,L$, are the hidden layer spaces.
@@
Next, we combine Definitions˜3.5 and 4.1.
Definition 4.2 (Deep integral RKBS).
Let $\mathcal{H}^{\rm deep}$ be an RKBS of depth $L$. If the RKBS $\mathcal{H}_{\ell}$ are integral for all $\ell=0,\dots,L$,
$\mathcal{H}_{\ell}=\{f_{\mu_{\ell}}:\mathcal{X}_{\ell}\to\mathcal{X}_{\ell+1}:\mu_{\ell}\in\mathcal{M}(\Theta_{\ell},\mathcal{X}_{\ell+1})\},$
@@
where $\rho_{\ell}(x,\cdot)\in\mathcal{C}_{0}(\Theta_{\ell})$ for all $x\in\mathcal{X}$. We call $\mathcal{H}^{\rm deep}$ an integral RKBS of depth $L$ and, if $L>1$, a deep integral RKBS.
@@
Thus, a function $f^{\rm deep}:\mathcal{X}\to\mathcal{Y}$ in a deep integral RKBS $\mathcal{H}^{\rm deep}$ of depth $L$ has the form
@@
$\begin{cases}x^{(0)}=x&\in\mathcal{X}\\[5.0pt] x^{(\ell+1)}=\displaystyle{\int_{\Theta_{\ell}}\rho_{\ell}(x^{(\ell)},\theta_{\ell}){\rm d}\mu_{\ell}(\theta_{\ell})}&\in\mathcal{X}_{\ell+1},\qquad\ell=0,\dots,L,\\ f^{\rm deep}(x)=x^{(L+1)}&\in\mathcal{Y}\end{cases}$ (8)
where $\mu_{\ell}\in\mathcal{M}(\Theta_{\ell},\mathcal{X}^{\ell+1})$ for any $\ell=0,\ldots,L$.